### v1

# Prompt-OT: An Optimal Transport Regularization Paradigm for Knowledge Preservation in Vision-Language Model Adaptation

Abstract
Vision-language models (VLMs) such as CLIP demonstrate strong performance but struggle when adapted to downstream tasks. Prompt learning has emerged as an efficient and effective strategy to adapt VLMs while preserving their pre-trained knowledge. However, existing methods still lead to overfitting and degrade zero-shot generalization. To address this challenge, we propose an optimal transport (OT)-guided prompt learning framework that mitigates forgetting by preserving the structural consistency of feature distributions between pre-trained and fine-tuned models. Unlike conventional point-wise constraints, OT naturally captures cross-instance relationships and expands the feasible parameter space for prompt tuning, allowing a better trade-off between adaptation and generalization. Our approach enforces joint constraints on both vision and text representations, ensuring a holistic feature alignment. Extensive experiments on benchmark datasets demonstrate that our simple yet effective method can outperform existing prompt learning strategies in base-to-novel generalization, cross-dataset evaluation, and domain generalization without additional augmentation or ensemble techniques. The code is available at https://github.com/ChongQingNoSubway/Prompt-OT
1 Introduction
Figure 1: (a) Comparison of point-wise constraints vs. our OT-based constraints: Unlike rigid point-wise alignment, our loss captures cross-instance relationships, effectively modeling correlations both within and between classes. (b) The error contours for base and novel tasks without constraint (top left), with point-wise constraint (bottom left), and our OT-based constraints (bottom right). Our OT-based constraint enlarges feasible parameter domains, striking a balance between adaptation and generalization.
Foundational vision-language models (VLMs), such as CLIP [19] and ALIGN [8], trained on large-scale datasets to align image-text pairs have demonstrated exceptional generalization capabilities across a variety of downstream tasks [11, 4, 15, 31]. However, recent studies have shown that VLMs face challenges in maintaining generalization when adapted to downstream tasks via e.g. fine-tuning [10], particularly for tasks with limited samples such as few-shot learning.
To tackle this issue, recent studies [29, 28] have shown that introducing learnable prompt tokens while keeping the original pre-trained weights fixed is an effective strategy for fine-tuning VLMs on downstream tasks (a.k.a. Prompt Learning). Following this trend, many efforts have been devoted to exploring different prompt learning strategies, either through the text branch alone [29, 28] or through both the text and image branches [9, 26]. However, early prompt learning methods [29, 28, 9] often lead to overfitting in task-specific datasets and sacrifice zero-shot generalization. To address this limitation, more recent studies  [22, 10] introduce explicit regularization to maintain the feature consistency between the finetuned and the frozen pre-trained (i.e. zero-shot) VLMs. These methods mainly employ a point/sample-wise metric (e.g. MSE/MAE) for the feature alignment between the zero-shot and finetuned models.
However, we argue that this may not be an optimal way to impose the aforementioned feature consistency due to the following reasons: (i) point-wise constraints may fail to capture cross-instance relationships, limiting their ability to account for the geometry of the underlying manifold and effectively model both within-class and cross-class relations. (ii) Point-wise constraints might be too restrictive111 A similar conclusion is also drawn in a well-known work [12] for related tasks. We have more discussion for this method in Sec. 2, as different features and samples have different importance. This may cause the features to learn from prompted VLM to be shrunk to a narrow searching space. This limitation can hinder both adaptation and generalizability.
To address these challenges, we propose an optimal transport (OT)-guided VLMs prompt tuning method, which leverages OT to mitigate knowledge forgetting between the embedding features of the zero-shot and the finetuned VLMs. The advantages of OT-guided prompts are multi-fold. First, OT naturally captures cross-instance relationships, since the transport map establishes an interactive coupling among instances by redistributing mass based on their mutual relationships. Second, we prove that OT expands the feasible domain for learnable prompt tokens. A broader feasible domain introduces more potential local minima, provides greater flexibility for the optimizer to explore, and allows easier hyperparameter adjustment, increasing the likelihood of achieving an optimal balance between adaptation and generalization. The conceptual illustration is shown in Fig. 1. Third, rather than independently constraining text and vision representations and merely computing a weighted sum of their losses, we incorporate the fundamental concept of an instance from CLIP, which inherently integrates both vision and text modalities. Consequently, our design enforces constraints on both modalities simultaneously for each instance. It is also worth noting that, despite its simple yet effective design, our method outperforms several recent competitive approaches without relying on additional components such as ensemble methods or data augmentation techniques. In summary, our work makes the following key contributions:
  • We propose an OT-based constraint scheme to mitigate knowledge forgetting, addressing the core limitations of previous constraint-based prompt learning methods.
  • Inspired by the fundamental concept of an instance in CLIP, we impose constraints on both modalities simultaneously for each instance.
  • Empirical evaluations across three popular benchmarks (i.e., Base-To-Novel Generalization, Cross-Dataset Evaluation, and Domain Generalization), along with ablation studies, demonstrate the effectiveness and superiority of our approach over recent alternative methods, such as PromptSRC [10], TCP [26], and QMaPLe [7].
2 Related Work
Adaption to Downstream tasks. Two popular ways to adapt a model for downstream tasks are full fine-tuning and linear probing. However, full fine-tuning can weaken generalization, while linear probing often yields poor performance, leading to both being problematic for tasks needing accuracy and generalization, like open-vocabulary image recognition and object detection [9]. To address this, researchers add extra learnable components without altering the original pre-trained weights, attempting to preserve both performance and generalization.[29]. One representative approach is known as prompt-based methods [29, 28, 30, 26]. In these methods, the pre-trained model is adapted by incorporating a small number of learnable prompt tokens into the original input. For example, CoOp [29], CoCoOp [28], ProdGrad [30], TCP [26] introduced a set of learnable continuous vectors into the text input. Although there are some attempts to only include prompt tokens on the image branch, more recent works prefer to perform prompt learning on both branches jointly, such as MaPLe [9] and PromptSRC [10]. Instead, adapter-based methods often add some extra layers to refine features generated by CLIP [27, 6, 25].
Consistency-based Learning. Instead of focusing on the design of the architecture, there are methods to enforce consistency between the trainable and pre-trained models to reduce overfitting and forgetting on downstream task fine-tuning. ProGrad [30] aligns gradient directions with the pre-trained model, while PromptSRC [10] imposes consistency on both embeddings and logits. A similar idea in [22] employs adapters, data augmentations, and LLM knowledge. Though PromptSRC is closest to our method, it only applies pairwise constraints that match average embedding distributions. We argue this is too restrictive and overlooks structural information, as discussed in Sec. 3.3. A related work [12] in continual learning attempts to solve severe restriction issues through a Fisher Information-based constraint, while it requires a substantial approximation during computing and needs the corresponding weights on the frozen model, which makes it infeasible to our case as the prompt tokens are only adding to the adapted model.
Optimal Transport for VLMs. Some recent studies also use OT to improve learning on CLIP, such as PLOT [1], Dude [16], and AWT [32]. PLOT employs multiple learnable text prompts and uses OT to align text and vision features. Dude leverages unbalanced OT to match class-specific and domain-shared text features (augmented by an LLM) with visual features. AWT, designed for zero-shot learning, applies various image transformations and LLM-augmented text prompts, then uses OT to measure the distance between an input image and candidate labels. Different from those methods, which typically use OT with augmented inputs to align vision and text features, our method applies OT as a regularization to align the vision-text joint distributions of the pre-trained and adapted models, without any need for augmentation.
Augmentation for VLMs. Several recent studies introduce extra augmentation to boost VLMs adaptation. For example, [10, 26] use more hand-crafted text templates beyond “a photo of a [category]” or design their own prompts, while [22, 32] leverage LLMs for diverse semantic prompts. Others [5] rely on Stable Diffusion [21] to generate multi-view images for better test-time tuning. These methods show that high-quality augmentation can significantly improve performance on downstream tasks. However, not everyone can access such data-augmentation resources. Thus, our method is designed to work without any extra augmentation, yet still delivers competitive results.
3 Method
Figure 2: The overview of our proposed framework. Only the prompt tokens are trainable, and the rest of the weights in both the zero-shot encoders and the adapted encoders are frozen. Despite the cross-entropy $\mathcal{L}_{ce}$ adopted, we also adopt the proposed joint optimal transport loss $\mathcal{L}_{\mathrm{jot}}$ between joint zero-short representation and adapted representation to constrain the model.
3.1 Preliminaries
We first briefly present the basic pipeline of CLIP as well as the principle of Prompt Learning.
Vision Encoding. An image $I$ is first divided into $M$ patches and then projected to latent space, which is denoted as $\bm{V}^{0}=\{\bm{v}_{1}^{0},\cdots,\bm{v}_{M}^{0}\}$. Afterward, a learnable class token $\bm{v}_{cls}^{0}$ is concatenated as the global representation, denoted as $[\bm{V}^{0},\bm{v}_{cls}^{0}]$. These tokens subsequently are fed into a sequence of transformer blocks:
$[\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (1)
where $l=1,\cdots,L_{I}$ and $L_{I}$ denote the number of vision transformer blocks in the image encoder of CLIP. Afterward, the final image representation in V-L space is obtained through an additional linear projection:
$\bm{h}=f_{\texttt{VisionProj}}(\bm{v}_{cls}^{L_{I}}).$ (2)
Text Encoding. Each text prompt is tokenized to $N$ tokens, denoted as $\bm{T}^{0}=\{\bm{t}_{1}^{0},\cdots,\bm{t}_{N}^{0}\}$. Likewise, these tokens then feed to text transformer blocks,
$\bm{T}^{l}$ $=\texttt{Trans}^{l}_{\texttt{Text}}(\bm{T}^{l-1}),$ (3)
where $l=1,\cdots,L_{T}$ and $L_{T}$ denote the number of transformer blocks in the text encoder of CLIP. Similarly, the final representation is obtained by the last token through a linear projection:
$\bm{g}=f_{\texttt{TextProj}}(\bm{t}_{N}^{L_{T}}).$ (4)
Prediction. In a standard setting of zero-shot classification [9], the text encoder generates the hand-crafted text prompt using the common template ”a photo of a [category]”. Given class labels $y\in\{1,\cdots,C\}$, we can obtain the text embedding for each class by filling out the template with the corresponding class name as $(\bm{g})_{y}$. As mentioned earlier, we also obtained the image representation for an image as $\bm{h}$. The prediction can be performed as:
$p(\bar{y}|I)=\frac{\text{exp}(sim(\bm{h},\bm{g}_{y})/\tau)}{\sum_{k=1}^{C}\text{exp}(sim(\bm{h},\bm{g}_{i}/\tau)},$ (5)
where $sim(\cdot,\cdot),\tau$ denote the cosine similarity and a pre-defined temperature, respectively.
Prompt Learning. Now, we will delve into how to perform prompt learning, which often appends some learnable tokens to text or vision branches as some transformer layers. Formally, we denote $\mathcal{S}_{\texttt{vision}}$ and $\mathcal{S}_{\texttt{text}}$ as the layers to integrate learnable prompt tokens for vision encoder and text encoder, respectively. We can denote the $K_{I}$ vision prompt tokens and $K_{T}$ text prompt tokens at $(l+1)$-th block as $\bm{P}^{l}=\{\bm{p}^{l}_{1},\cdots,\bm{p}^{l}_{K_{I}}\}$ and $\bm{Q}^{l}=\{\bm{q}^{l}_{1},\cdots,\bm{q}^{l}_{K_{T}}\}$, respectively. The prompting for vision encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{vision}}$) can be then presented as:
[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{P}^{l-1},\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (6)
and for text encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{text}}$) can be then presented as:
[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{T}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Text}}([\bm{Q}^{l-1},\bm{T}^{l-1}]).$ (7)
It is worth mentioning that this is known as Independent Vision-Language Prompting (IVLP), if $\bm{P}^{l}$ and $\bm{Q}^{l}$ are defined independently. This is used in [20, 10], and we will also use this baseline in our framework, following [10]. MaPLe [9] imposes an additional linear layer to share the knowledge between $\bm{P}$ and $\bm{Q}$. In order to adapt classification on a downstream dataset, we can simply update the prompt tokens via cross-entropy loss $\mathcal{L}_{CE}$ between predictions (Eq. 5) and ground truth labels.
3.2 Vision-Text Joint Optimal Transport
In this section, we will delve into the design of our framework. During training, this framework involves a CLIP with learnable prompts, i.e., the adapted model and a frozen pre-trained CLIP, while during inference, we only need the adapted CLIP. The overview of this framework is shown in Fig. 2.
A key principle in CLIP is that an instance has representations from both the text domain and the image domain. Therefore, we should constrain them simultaneously for each instance. Suppose we have a dataset $\mathcal{D}=\{(I_{i},y_{i})\}_{i=1}^{n}$, where $I_{i},y_{i}$, and $n$ denote the $i$th input image, its corresponding label, and number of samples, respectively. For each instance $i\in[n]$, the vision representation is obtained by the vision encoder as $\bm{h}^{i}$. Since the real text caption for the image is often unknown, we use the standard template with its category as the proxy, and accordingly, we can obtain the text representation as $\bm{g}^{i}$. For simplicity, we concatenate both representations as the joint representation of an instance $i$, denoted as $\bm{x}_{i}=\texttt{concat}(\bm{h}^{i},\bm{g}^{i})$. Now, we have the distribution of joint representation of the dataset $\mathcal{D}$ as $\mathcal{X}=\{\bm{x}^{i}\}_{i=1}^{n}$ by the adapted model. We also need to obtain the joint representation distribution using an additional CLIP model, i.e., the CLIP model with its original pre-trained weights, while keeping all parameters frozen. We denote it as $\bm{X}_{\texttt{zs}}=\{\bm{x}_{\texttt{zs}}^{i}\}_{i=1}^{n}$. It is worth mentioning that $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{i}$ are obtained from the same input instance $(I_{i},y_{i})$, while with different CLIP models (i.e. one keep updating and one frozen).
Our proposed regularization term is presented as: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the proposed constraint lies in Optimal Transport [23] is defined as:
$\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})=\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),$ (8)
where $c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})$ denotes a cost function (e.g. linear and quadratic cost) between $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{j}$ $\bm{\gamma}$ denotes a transport map, $\mathcal{U}(\bm{a},\bm{b})$ denotes set of all valid transport maps,
$\mathcal{U}(\bm{a},\bm{b})$ =$=\left\{\bm{\gamma}\in\mathbb{R}^{n\times n}_{+}\ \middle|\ \bm{\gamma}\mathbf{1}_{n}=\bm{a},\ \bm{\gamma}^{\top}\mathbf{1}_{n}=\bm{b}\right\},$ (9)
$\bm{a}$ $=\frac{1}{n}\mathbf{1}_{n},\quad\bm{b}=\frac{1}{n}\mathbf{1}_{n},$
and $\mathbf{1}_{n}$ is an $n$-dimensional vector of ones.
In summary, this proposed loss aims to preserve the consistency between updated joint representation and zero-shot representation while offering a larger search feasible domain of prompts and taking cross-instance from the same class or different class into consideration. Below, we will provide a detailed analysis with theoretical justification.
3.3 Theoretical Analysis
Modeling cross-instance correlations can potentially characterize both within-class relations and between-class relations. Under the proposed OT-based constraint, each element in the transport map $\bm{\gamma}$ represents the amount of mass transported between two instances from the zero-shot representation and adapted representation. The learned transport map inherently captures cross-instance correlations because OT enforces a structured assignment, where similar instances are mapped with lower cost while maintaining the geometry of the data distribution. This process happens automatically because OT optimizes the movement of probability mass, ensuring that within-class relations remain compact while between-class relations reflect meaningful transitions, ultimately shaping a well-organized feature space.
Here, we focus more on how using OT-loss can potentially enlarge the feasible parameter space, which potentially results in easier optimization and better convergence.
We first give a general definition of the loss of point-wise constraint: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the point-wise constraint is defined as:
$\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})=\frac{1}{n}\sum_{i=1}^{n}c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{i}).$ (10)
We observe that it can be presented in a similar form of optimal transport: {lemma} An equivalent form of Definition 3.3 is presented as:
$\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})=\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),$ (11)
where $\forall i\neq j,\hat{\gamma}_{ij}=0$ and $\forall i=j,\hat{\gamma}_{ij}=1/n$.
{lemma}
Let $\bm{X}_{\texttt{zs}}$ be a zero-shot representation distribution and $\epsilon>0$ be a tolerance of the constraint. Suppose there exists a set $\mathcal{X}_{\texttt{pw}}$ such that for all $\bm{X}\in\mathcal{X}_{\texttt{pw}}$, $\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}$. Similarly, let there exist a set $\mathcal{X}_{\texttt{ot}}$ such that for all $\bm{X}\in\mathcal{X}_{\text{ot}}$, $\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}$. Then, it must hold that $\mathcal{X}_{\texttt{pw}}\subseteq\mathcal{X}_{\texttt{ot}}$.
Proof.
We first note that, clearly, according to Definition 3.2, $\hat{\gamma}\in\mathcal{U}(\bm{a},\bm{b})$ in Lemma 3.3 is a valid transport map. Accordingly,
$\underbrace{\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})}_{\mathcal{L}_{\mathrm{pw}}}\geq\underbrace{\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})}_{\mathcal{L}_{\mathrm{jot}}}.$ (12)
This is due to $\mathcal{L}_{\mathrm{jot}}$ minimizes over all valid transport plans while $\mathcal{L}_{\mathrm{pw}}$ is applying one certain valid transport plans that not necessarily is the optimal one. Therefore, given $\bm{X}$, if $\mathcal{L}_{\text{pw}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}$ holds, $\mathcal{L}_{\text{jot}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}$ must hold. Hence, any $\bm{X}\in\mathcal{X}_{pw}$ satisfies the condition for $\mathcal{X}_{ot}$. ∎
Now, we will link this theorem to the feasible parameter space of the learnable prompt tokens. First, we recap the objective function with consistency constraints:
$\min_{\mathcal{P},\mathcal{Q}}\mathcal{L}_{ce}+\lambda R(\bm{X},\bm{X}_{\texttt{zs}}),$ (13)
where $\mathcal{L}_{ce}$ denotes the cross-entropy loss widely used in adaptation for CLIP. $\mathcal{P}=\{\bm{P}^{l}\}_{l\in\mathcal{S}_{\texttt{vision}}},\mathcal{Q}=\{\bm{Q}^{l}\}_{l\in\mathcal{S}_{\texttt{text}}}$ denote all learnable tokens from different layers. $\lambda\geq 0$ is a tuning hyper-parameter.
This objective can be considered as the generalized Lagrange function of the following optimization problem:
$\min_{\mathcal{P},\mathcal{Q}}\quad\mathcal{L}_{ce}$ (14)
$s.t.~{}R(\bm{X},\bm{X}_{\texttt{zs}})-\epsilon^{2}\leq 0,$
where $\epsilon^{2}$ is a certain constant. Now, we have the following conclusion: {theorem} Let any $\Omega_{\texttt{pw}}$ be a feasible domain of $(\mathcal{P},\mathcal{Q})$ to the optimization problem defined in Eq. 14 under the point-wise constraint $\mathcal{L}_{\mathrm{pw}}$ (i.e., with $R:=\mathcal{L}_{\mathrm{pw}}$). Similarly, let any $\Omega_{\texttt{ot}}$ be a feasible domain of $(\mathcal{P},\mathcal{Q})$ to Eq. 14 under the optimal transport constraint $\mathcal{L}_{\mathrm{jot}}$ (i.e., with $R:=\mathcal{L}_{\mathrm{jot}}$). Then, the following inclusion holds:
$\Omega_{\texttt{pw}}\subseteq\Omega_{\texttt{jot}}.$ (15)
Proof.
We define the functions $g_{\mathrm{pw}}:\Omega_{\texttt{pw}}\rightarrow\mathcal{X}_{\texttt{pw}}^{*}$ and $g_{\mathrm{jot}}:\Omega_{\texttt{ot}}\rightarrow\mathcal{X}_{\texttt{ot}}^{*}$ to represent the forward propagation of the model with dataset $\mathcal{D}$ under the point-wise constraint and the optimal transport constraint, respectively.
Now, consider any $(\mathcal{P},\mathcal{Q})\in\Omega_{\texttt{pw}}$. Its representation distribution $\bm{X}=g_{\mathrm{pw}}(\mathcal{P},\mathcal{Q};\mathcal{D})$ must satisfy
$\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}.$
According to Lemma 3.3, any $\bm{X}\in\mathcal{X}_{\texttt{pw}}^{*}$ also satisfies
$\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})\leq\epsilon^{2}.$
This implies that any $\bm{X}\in\mathcal{X}_{\texttt{pw}}^{*}$ also belongs to the set $\mathcal{X}_{\texttt{ot}}^{*}$.
Since for any $\bm{X}\in\mathcal{X}_{\texttt{ot}}^{*}$, we can find at least one corresponding $(\mathcal{P},\mathcal{Q})\in\Omega_{\texttt{ot}}$, it follows that $\Omega_{\texttt{pw}}\subseteq\Omega_{\texttt{jot}}$. ∎
{remark}
Theorem 3.3 immediately suggests that using optimal transport constraints allows a broader feasible parameter domain of learnable tokens $(\mathcal{P},\mathcal{Q})$. A larger feasible domain often indicates there are potential more local minimal existing and provides more pathways for the optimizer to explore, which increases the chances of finding a better solution and alleviates a heavy hyperparameter search, i.e., a uniform hyperparameter can achieve reasonable performance across different datasets.
3.4 Training and Inference
Training. Now we assemble the proposed loss in training, which is presented as
$\min\mathcal{L}_{ce}+\lambda\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}}).$ (16)
we employ mini-batch OT to take the mini-batch training nature, which involves computing OT distance within each batch. It is noteworthy that the mini-batch is well-bounded as studied by multiple literatures, including [3, 17, 18].
Inference. Inference does not require any computation of OT and does not need the zero-shot CLIP model. Only the adapted CLIP model is used.
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bolded. HM indicates the harmonic mean. Prompt-OT represents the best accuracy achieved across different datasets by searching over epochs. It is not intended for direct comparison but is included here for reference.
(a) Average
Base Novel HM
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
Co-CoOp 80.47 71.69 75.83
ProGrad 82.48 70.75 76.16
KgCoOp 80.73 73.60 77.00
MaPLe 82.28 75.14 78.55
PromptSRC 84.26 76.10 79.97
ProDA 81.56 72.30 76.65
PLOT 83.98 71.72 77.37
TCP 84.13 75.36 79.51
QMaPLe 83.02 75.57 79.12
Prompt-OT 84.81 76.25 80.30
Prompt-OT 84.82 76.63 80.52
(b) ImageNet
Base Novel HM
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
Co-CoOp 75.98 70.43 73.10
ProGrad 77.02 66.66 71.46
KgCoOp 75.83 69.96 72.78
MaPLe 76.66 70.54 73.47
PromptSRC 77.60 70.73 74.01
ProDA 75.40 70.23 72.72
PLOT 77.30 69.87 73.40
TCP 77.27 69.87 73.38
QMaPLe 76.93 70.73 73.70
Prompt-OT 77.90 69.83 73.65
Prompt-OT 77.63 70.13 73.69
(c) Caltech101
Base Novel HM
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
Co-CoOp 97.96 93.81 95.84
ProGrad 98.02 93.89 95.91
KgCoOp 97.72 94.39 96.03
MaPLe 97.74 94.36 96.02
PromptSRC 98.10 94.03 96.02
ProDA 98.27 93.23 95.68
PLOT 98.53 92.80 95.58
TCP 98.23 94.67 96.42
QMaPLe 97.97 95.00 96.46
Prompt-OT 98.37 94.50 96.39
Prompt-OT 98.33 94.90 96.59
(d) OxfordPets
Base Novel HM
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
Co-CoOp 95.20 97.69 96.43
ProGrad 95.07 97.63 96.33
KgCoOp 94.65 97.76 96.18
MaPLe 95.43 97.76 96.58
PromptSRC 95.33 97.30 96.30
ProDA 95.43 97.83 96.62
PLOT 94.50 96.83 95.65
TCP 94.67 97.20 95.92
QMaPLe 95.67 97.63 96.64
Prompt-OT 95.50 97.03 96.26
Prompt-OT 95.70 97.23 96.46
(e) StanfordCars
Base Novel HM
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
Co-CoOp 70.49 73.59 72.01
ProGrad 77.68 68.63 72.88
KgCoOp 71.76 75.04 73.36
MaPLe 72.94 74.00 73.47
PromptSRC 78.27 74.97 76.58
ProDA 74.70 71.20 72.91
PLOT 79.07 74.80 76.88
TCP 80.80 74.13 77.32
QMaPLe 75.00 73.67 74.33
Prompt-OT 80.60 74.80 77.59
Prompt-OT 81.17 74.70 77.80
(f) Flowers102
Base Novel HM
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
Co-CoOp 94.87 71.75 81.71
ProGrad 95.54 71.87 82.03
KgCoOp 95.00 74.73 83.65
MaPLe 95.92 72.46 82.56
PromptSRC 98.07 76.50 85.95
ProDA 97.70 68.68 80.66
PLOT 97.93 73.53 83.99
TCP 97.73 75.57 85.23
QMaPLe 96.43 74.33 83.95
Prompt-OT 98.17 77.03 86.32
Prompt-OT 98.23 77.53 86.66
(g) Food101
Base Novel HM
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
Co-CoOp 90.70 91.29 90.99
ProGrad 90.37 89.59 89.98
KgCoOp 90.50 91.70 91.09
MaPLe 90.71 92.05 91.38
PromptSRC 90.67 91.53 91.10
ProDA 90.30 88.57 89.43
PLOT 89.80 91.37 90.58
TCP 90.57 91.37 90.97
QMaPLe 90.63 92.10 91.36
Prompt-OT 90.67 91.63 91.15
Prompt-OT 90.70 92.00 91.35
(h) FGVCAircraft
Base Novel HM
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
Co-CoOp 33.41 23.71 27.74
ProGrad 40.54 27.57 32.82
KgCoOp 36.21 33.55 34.83
MaPLe 37.44 35.61 36.50
PromptSRC 42.73 37.87 40.15
ProDA 36.90 34.13 35.46
PLOT 42.13 33.73 37.46
TCP 41.97 34.43 37.83
QMaPLe 39.10 34.90 36.88
Prompt-OT 44.47 36.57 40.12
Prompt-OT 44.60 36.67 40.24
(i) SUN397
Base Novel HM
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
Co-CoOp 79.74 76.86 78.27
ProGrad 81.26 74.17 77.55
KgCoOp 80.29 76.53 78.36
MaPLe 80.82 78.70 79.75
PromptSRC 82.67 78.47 80.52
ProDA 78.67 76.93 77.79
PLOT 82.20 73.63 77.68
TCP 82.63 78.20 80.35
QMaPLe 81.33 78.27 79.77
Prompt-OT 82.53 78.90 80.68
Prompt-OT 82.67 78.90 80.74
(j) DTD
Base Novel HM
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
Co-CoOp 77.01 56.00 64.85
ProGrad 77.35 52.35 62.45
KgCoOp 77.55 54.99 64.35
MaPLe 80.36 59.18 68.16
PromptSRC 83.37 62.97 71.75
ProDA 80.67 56.48 66.44
PLOT 81.97 43.80 57.09
TCP 82.77 58.07 68.25
QMaPLe 80.77 57.63 67.27
Prompt-OT 83.63 64.00 72.51
Prompt-OT 83.5 64.27 72.63
(k) EuroSAT
Base Novel HM
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
Co-CoOp 87.49 60.04 71.21
ProGrad 90.11 60.89 72.67
KgCoOp 85.64 64.34 73.48
MaPLe 94.07 73.23 82.35
PromptSRC 92.90 73.90 82.32
ProDA 83.90 66.00 73.88
PLOT 93.70 62.67 75.11
TCP 91.63 74.73 82.32
QMaPLe 94.30 79.47 86.25
Prompt-OT 93.10 75.27 83.21
Prompt-OT 92.77 77.10 84.19
(l) UCF101
Base Novel HM
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
Co-CoOp 82.33 73.45 77.64
ProGrad 84.33 74.94 79.35
KgCoOp 82.89 76.67 79.65
MaPLe 83.00 78.66 80.77
PromptSRC 87.10 78.80 82.74
ProDA 85.23 71.97 78.04
PLOT 86.60 75.90 80.90
TCP 87.13 80.77 83.83
QMaPLe 85.10 77.50 81.12
Prompt-OT 87.93 79.23 83.36
Prompt-OT 87.73 79.50 83.41
4 Experiment
4.1 Setups
We follow the experimental setups of [29, 9, 26] and evaluate on three tasks: Base-to-Novel Generalization, Cross-Data Evaluation, and Domain Generalization. For Base-to-Novel Generalization, we train on base classes and evaluate on both base and novel classes. Following [10], we use 11 datasets. To ensure fairness, we adopt two strategies: using the same epoch for all datasets (adhering to prior methods) and searching for the best epoch on each dataset (as a reference, only to show our method’s full potential). For Cross-Data Evaluation, we train on ImageNet [2] and test on 10 other datasets without extra fine-tuning. For Domain Generalization, we again train on ImageNet [2] and evaluate on four out-of-distribution datasets.
Table 2: Performance of PromptOT on cross-dataset evaluation and its comparison to existing methods. Here, the model is trained on the ImageNet dataset and evaluated on ten other datasets in a zero-shot setting.
Caltech Pets Cars Flowers Food Aircraft SUN397 DTD EuroSAT UCF Ave.
CoOp 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
Co-CoOp 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
MaPLe 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
PromptSRC 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
PLOT 92.07 90.10 65.70 69.23 86.23 25.00 61.67 38.60 47.83 67.00 64.34
TCP 93.97 91.25 64.69 71.21 86.69 23.45 67.12 44.35 51.45 68.73 66.29
QCoOp 94.07 90.53 65.97 71.33 86.23 22.73 66.80 44.20 48.23 69.17 65.93
Prompt-OT 94.03 90.47 65.87 71.27 86.43 23.63 67.20 46.67 50.57 69.03 66.52
Baselines. Despite compared with original CLIP [19], we select multiple recent prompt learning methods for comparison, including: CoOp [29], CoCoOp [28], ProDA [14], MaPLe [9], ProGrad [10], PromptSRC [10], PLOT [1], TCP [26], and QMaPLe/QCoOP [7]. It is worth mentioning that a portion of the methods are only designed for partial tasks, while our method does not have any obstacles to implementation for all tasks. For the sake of a fair comparison and proof of concept, we will skip comparing methods that introduce substantial additional knowledge (e.g., LLMs or larger CLIP models for knowledge distillation) in our main experiments, such as those proposed in [22, 13, 32, 24].
Implementation details. For fair comparison, We use a ViT-B/16-based CLIP model, consistent with prior work [9, 10, 26]. Prompts are initialized from a normal distribution, except for the text prompts in the text input, which begin with the embedding of ”a photo of”. We fix the learning rate at 0.005 and set $\lambda=10$ for all benchmarks. All reported results are the average of three runs, without any extra augmentation. Baseline results are taken from their respective papers. For more detail, refer to Appendix A.
4.2 Main Results
Base-to-Novel Generalization. As shown in Table 1, our method significantly outperforms zero-shot CLIP, confirming its effectiveness. In particular, we achieve gains of 15.47%, 2.03%, and 8.60% in base task accuracy, novel task accuracy, and their harmonic mean, respectively. More importantly, when averaged over 11 datasets, our approach surpasses all recent baselines. Specifically, it achieves 84.81% (base task accuracy), 76.26% (novel task accuracy), and 80.30% (harmonic mean), which is an improvement of 0.55%, 0.15%, and 0.33% over the previous state-of-the-art, PromptSRC.
Cross-Dataset Evaluation. Table 2 shows that our method attains an average accuracy of 66.52% across 11 datasets, outperforming all baselines and exceeding PromptSRC by 0.62%.
Domain Generalization. As shown in Table 3, we observe that our method achieves the highest average accuracy on four out-of-distribution datasets.
Table 3: Performance on domain generalization. These baseline models are trained on ImageNet and evaluated on four out-of-distribution ImageNet datasets.
ImNetV2 ImNetS ImNetA ImNetR Ave.
CLIP 60.83 46.15 47.77 73.96 57.17
UPT 64.35 48.66 50.66 76.24 59.98
CoOp 64.20 47.99 49.71 75.21 59.28
Co-CoOp 64.07 48.75 50.63 76.18 59.90
ProGrad 64.73 47.61 49.39 74.58 59.07
KgCoOp 64.10 48.97 50.69 76.70 60.11
MaPLe 64.07 49.15 50.90 76.98 60.26
PromptSRC 64.35 49.55 50.90 77.80 60.65
TCP 64.60 49.50 51.20 76.73 60.51
QCoOp 63.87 48.93 51.10 76.90 60.20
Prompt-OT 64.35 49.40 51.63 77.40 60.70
Our method demonstrates robust performance across three key evaluations. In the base-to-novel generalization setup, it substantially outperforms zero-shot CLIP and surpasses the previous state-of-the-art in base accuracy, novel accuracy, and harmonic mean. In cross-dataset evaluation, it achieves the highest average accuracy over 11 diverse datasets, again surpassing all baselines. Finally, in domain generalization, our method obtains the best overall accuracy without relying on prompt ensembles or text augmentation, underscoring both its effectiveness and simplicity.
5 Analysis
In this section, we will delve into a more detailed analysis of our proposed method. All experiments will be conducted consistently with the setups of base-to-novel generalization. For fairness, we exclude all other components, and only apply regularization on the IVLP.
Ours VS Point-wise Constraints. To illustrate the superiority of the proposed method over point-wise constraints, we evaluate it against several baselines: (i) naive $L_{2}$ constraints between $\bm{X}$ and $\bm{X}_{\texttt{zs}}$ with the same regularization strength $\lambda$, (ii) Adaptor-Cos [22], which introduces adapters at the end of the adapted encoders and employs cosine similarity as the point-wise constraint between representations, (iii) SRC [10], which applies $L_{1}$ losses for vision and text representations separately and a KL-divergence between prediction probabilities for each instance. (iv) Separate OT (see Appendix A), which applies OT-loss for vision and text representations separately. For a fair comparison, we exclude all additional augmentations, such as ensemble techniques and LLM-based augmentation. Our primary observation is that all these methods achieve a better balance between novel and base tasks than the model without any constraints, with a minimum improvement of 1.83% in harmonic mean accuracy. More importantly, the OT-based methods outperform all recent approaches based on point-wise constraints. Even without joint representation, applying OT for two domains separately can outperform all other point-wise methods. Additionally, compared to SRC, the previous state-of-the-art method, we achieved gains of 0.4%, 1.35%, and 0.93% in base task accuracy, novel task accuracy, and harmonic mean accuracy, respectively. Compared to $L_{2}$ constraints, we observe that with the same regularization strength $\lambda$, $L_{2}$ constraints, while preserving good generalizability to novel tasks, fail to yield any improvements on the base task. This suggests that $L_{2}$ constraints may be overly restrictive for model adaptation. In contrast, our proposed approach effectively balances performance across both the base and novel tasks, demonstrating that modeling cross-instance correlation while allowing the model greater flexibility in updating leads to an optimal trade-off, as expected in Fig. 1.
Table 4: Comparison of performance with point-wise constraints.
Reg. Method Base Novel HM
w/o Constraints 84.20 71.21 77.16
Point-wise $L_{2}$ between $\bm{X},\bm{X}_{\texttt{zs}}$ 84.20 74.68 79.16
Adpator-Cos
From CoPormpt [22]
83.12 75.25 78.99
SRC
From PromptSRC [10]
84.41 74.89 79.37
Separate OT 84.50 75.48 79.74
Our 84.81 76.25 80.30
Effectiveness of Joint Representation. Here, we investigate alternative approaches to imposing constraints between zero-shot and adapted representations. Specifically, we examine four different strategies: (i) imposing the constraint solely on the text representation, (ii) imposing the constraint solely on the vision representation, (iii) applying constraints separately to text and vision representation distributions (i.e., Separate OT), and (iv) our proposed joint representation, which considers an instance includes both vision and text representations. As shown in Table 5, applying constraints to both modalities yields a significantly greater improvement (2.58% and 3.14% in HM) compared to imposing the constraint on a single modality (0.62% and 0.96% in HM). Moreover, our proposed joint representation obtain more gain than applying constraints on vision and text distributions separately. Our method leads to the highest increase in HM (3.14%), suggesting that it provides the most effective balance between minimizing errors in both novel and base classes.
Table 5: Ablation study on different modalities for imposing constraints.
Text Vision Scheme Base Novel HM
- 84.20 71.21 77.16
- 84.92 71.75 77.78
- 84.12 72.94 78.12
Separate 84.50 75.48 79.74
Joint 84.81 76.25 80.30
Sensitivity of $\lambda$. We further conduct a study to evaluate the impact of different values of $\lambda$ on adaptation. As shown in Fig. 3, as $\lambda$ increases, the constraints become stronger, leading to a decline in adaptation to the base class. Meanwhile, the accuracy on the novel class initially increases before decreasing, indicating that the zero-shot weights may not be optimal for both novel and base tasks. This suggests that some relaxation is necessary to achieve a well-balanced trade-off. Excessively strong constraints may prevent the model from reaching the low-error regions of both tasks in the optimization landscape. Empirically, we find that $\lambda=10$ achieves a good balance between generalization and adaptation.
Figure 3: The effectiveness of $\lambda$ on base-to-novel generalization tasks averaged over 11 datasets.
6 Conclusion
In this work, we proposed an Optimal Transport (OT) Regularization framework for vision-language prompt learning to mitigate forgetting and enhance adaptability. Unlike rigid point-wise constraints, our method uses OT to model cross-instance relationships and preserve the embedding structure between pre-trained and fine-tuned models. Theoretical analysis shows that OT constraints offer a broader feasible parameter space, enabling more robust prompt learning. Extensive experiments demonstrate that our approach outperforms state-of-the-art methods across various generalization settings. Future directions include incorporating adaptive regularization strength to improve trade-offs dynamically and exploring efficient transport solvers for broader applications.
Appendix A Implementation Detail
We train PromptOT for 30 epochs for the Base-to-Novel Generalization benchmark and 2 epochs for the remaining three benchmark settings, respectively. The respective epochs are fixed across all datasets.
Following a deep prompting strategy, we apply prompts to all transformer blocks for the first benchmark and the first three transformer blocks for the remaining benchmarks, and we use four text prompts and four vision prompts.
Following previous works [10], in For Base-to-Novel Generalization benchmark, we set the first half of classes as base classes and the remaining classes as novel classes. Models are only trained on base classes. The accuracy is reported based on the evaluation of base classes and novel classes, respectively.
A.1 Separate OT
The way of applying separate OT is similar to SRC loss. Suppose the adapted vision features are presented as $\bm{H}=\{\bm{h}^{i}\}_{i=1}^{n}$ and adapted text features are presented as $\bm{G}=\{\bm{g}^{i}\}_{i=1}^{C}$. Again, $n$ and $C$ denote the number of samples and number of classes, respectively. Similarly, we denote the zero-shot vision features and text features as $\bm{H}_{zs}$ and $\bm{H}_{zs}$, respectively. The Separate OT is denoted as:
$\mathcal{L}_{\texttt{SOT}}=\mathcal{L}_{\texttt{OT}}(\bm{H},\bm{H}_{zs})+\mathcal{L}_{\texttt{OT}}(\bm{G},\bm{G}_{zs}).$ (17)
A.2 Reproducibility
The code and model weights will be publicly released upon acceptance. In the supplementary zip file, we provide the key code for the configuration and implementation of our proposed method.
Our code is developed based on https://github.com/muzairkhattak/PromptSRC.

### v2

@@
 Our approach enforces joint constraints on both vision and text representations, ensuring a holistic feature alignment.
-Extensive experiments on benchmark datasets demonstrate that our simple yet effective method can outperform existing prompt learning strategies in base-to-novel generalization, cross-dataset evaluation, and domain generalization without additional augmentation or ensemble techniques.
+Extensive experiments on benchmark datasets demonstrate that our simple yet effective method outperforms existing prompt learning strategies in base-to-novel generalization, cross-dataset evaluation, and domain generalization, without requiring additional augmentation or ensemble techniques.
 The code is available at https://github.com/ChongQingNoSubway/Prompt-OT
 1 Introduction
+Foundational vision-language models (VLMs), such as CLIP [30] and ALIGN [15], trained on large-scale datasets to align image-text pairs, have demonstrated exceptional generalization capabilities across a variety of downstream tasks [18, 8, 23].
+However, recent studies have shown that VLMs face challenges in maintaining generalization when adapted to downstream tasks via e.g.
+fine-tuning [17], particularly for tasks with limited samples such as few-shot learning.
 Figure 1: (a) Comparison of point-wise constraints vs.
@@
 Our OT-based constraint enlarges feasible parameter domains, striking a balance between adaptation and generalization.
-Foundational vision-language models (VLMs), such as CLIP [19] and ALIGN [8], trained on large-scale datasets to align image-text pairs have demonstrated exceptional generalization capabilities across a variety of downstream tasks [11, 4, 15, 31].
-However, recent studies have shown that VLMs face challenges in maintaining generalization when adapted to downstream tasks via e.g.
-fine-tuning [10], particularly for tasks with limited samples such as few-shot learning.
 To tackle this issue, recent studies [44, 43] have shown that introducing learnable prompt tokens while keeping the original pre-trained weights fixed is an effective strategy for fine-tuning VLMs on downstream tasks (a.k.a.
@@
 MSE/MAE) for the feature alignment between the zero-shot and finetuned models.
-However, we argue that this may not be an optimal way to impose the aforementioned feature consistency due to the following reasons: (i) point-wise constraints may fail to capture cross-instance relationships, limiting their ability to account for the geometry of the underlying manifold and effectively model both within-class and cross-class relations.
-(ii) Point-wise constraints might be too restrictive111 A similar conclusion is also drawn in a well-known work [12] for related tasks.
+Despite this, we argue that this may not be an optimal way to impose the aforementioned feature consistency due to the following reasons: (i) Point-wise constraints may fail to capture cross-instance relationships, limiting their ability to account for the geometry of the underlying manifold and effectively model both within-class and cross-class relations.
+(ii) Point-wise constraints might be too restrictive [^1] A similar conclusion is also drawn in a well-known work [19] for related tasks.
 We have more discussion for this method in Sec.
-2, as different features and samples have different importance.
+2., as different features and samples have different importance.
 This may cause the features to learn from prompted VLM to be shrunk to a narrow searching space.
@@
 1.
-Third, rather than independently constraining text and vision representations and merely computing a weighted sum of their losses, we incorporate the fundamental concept of an instance from CLIP, which inherently integrates both vision and text modalities.
+Third, existing methods often constrain vision and text representations independently and then combine them with a weighted sum of losses.
+However, the vision encoder outputs features of size $B\times d$ (where $B$ and $d$ denote the batch size and the latent dimension, respectively), while the text encoder outputs features of size $C\times d$ (where $C$ denotes the number of classes).
+This mismatch can lead to imbalanced constraint strength across modalities.
+To address this, we build on the fundamental concept of an instance in CLIP, which inherently couples both vision and text.
+Our design therefore enforces constraints on vision and text simultaneously at the instance level, ensuring more balanced and structurally consistent multimodal representations.
 Consequently, our design enforces constraints on both modalities simultaneously for each instance.
@@
   • We propose an OT-based constraint scheme to mitigate knowledge forgetting, addressing the core limitations of previous constraint-based prompt learning methods.
-  • Inspired by the fundamental concept of an instance in CLIP, we impose constraints on both modalities simultaneously for each instance.
-  • Empirical evaluations across three popular benchmarks (i.e., Base-To-Novel Generalization, Cross-Dataset Evaluation, and Domain Generalization), along with ablation studies, demonstrate the effectiveness and superiority of our approach over recent alternative methods, such as PromptSRC [10], TCP [26], and QMaPLe [7].
+Beyond the conceptual design, we provide a theoretical justification showing that OT enlarges the feasible parameter space of learnable prompts, thereby enabling a more flexible and principled trade-off between adaptation and generalization.
+  • We propose to impose the OT constraint jointly on vision and text features of each instance.
+This balances the constraint strength across modalities and preserves structural consistency in multimodal representations, addressing the limitations of treating modalities independently.
+  • Empirical analysis (Table 4) highlights that Prompt-OT consistently outperforms all point-wise constraint methods with statistical significance.
+Across three popular benchmarks (Base-to-Novel Generalization, Cross-Dataset Evaluation, and Domain Generalization), our method achieves competitive or superior performance compared to recent approaches, including PLOT [2], PromptSRC [17], TCP [41], and QMaPLe [11], without relying on prompt ensembles or extra augmentation.
 2 Related Work
-Adaption to Downstream tasks.
+Adaptation to Downstream tasks.
 Two popular ways to adapt a model for downstream tasks are full fine-tuning and linear probing.
 However, full fine-tuning can weaken generalization, while linear probing often yields poor performance, leading to both being problematic for tasks needing accuracy and generalization, like open-vocabulary image recognition and object detection [16].
-To address this, researchers add extra learnable components without altering the original pre-trained weights, attempting to preserve both performance and generalization.[29].
+To address this, researchers add extra learnable components without altering the original pre-trained weights, attempting to preserve both performance and generalization [44].
 One representative approach is known as prompt-based methods [44, 43, 45, 41].
@@
 Consistency-based Learning.
-Instead of focusing on the design of the architecture, there are methods to enforce consistency between the trainable and pre-trained models to reduce overfitting and forgetting on downstream task fine-tuning.
+Instead of focusing on the design of the architecture, there are methods to enforce consistency between the trainable and pre-trained models, thereby reducing overfitting and forgetting during downstream task fine-tuning.
 ProGrad [45] aligns gradient directions with the pre-trained model, while PromptSRC [17] imposes consistency on both embeddings and logits.
@@
 3.3.
-A related work [12] in continual learning attempts to solve severe restriction issues through a Fisher Information-based constraint, while it requires a substantial approximation during computing and needs the corresponding weights on the frozen model, which makes it infeasible to our case as the prompt tokens are only adding to the adapted model.
+A related work [19] in continual learning attempts to solve severe restriction issues through a Fisher Information-based constraint, while it requires a substantial approximation during computing and needs the corresponding weights on the frozen model, which makes it infeasible for our case, as the prompt tokens are only added to the adapted model.
+Additionally, previous methods typically involve terms that constrain each modality independently.
+However, during CLIP training, the feature spaces of the two modalities are inherently imbalanced: the image encoder produces a $B\times d$ feature map, while the text encoder outputs a fixed $C\times d$ representation.
+This discrepancy leads to uneven constraint strength across modalities.
+To address this, we propose imposing constraints directly on the joint representation of each instance, ensuring better alignment.
+A detailed analysis is provided in Sec.
+5 and Table 5.
 Optimal Transport for VLMs.
@@
 Augmentation for VLMs.
-Several recent studies introduce extra augmentation to boost VLMs adaptation.
-For example, [10, 26] use more hand-crafted text templates beyond “a photo of a [category]” or design their own prompts, while [22, 32] leverage LLMs for diverse semantic prompts.
+Several recent studies introduce extra augmentation to boost VLMs’ adaptation.
+For example, [17, 41] use more hand-crafted text templates beyond “a photo of a [category]” or design their prompts, while [34, 46] leverage LLMs for diverse semantic prompts.
 Others [9] rely on Stable Diffusion [33] to generate multi-view images for better test-time tuning.
@@
 3 Method
-Figure 2: The overview of our proposed framework.
-Only the prompt tokens are trainable, and the rest of the weights in both the zero-shot encoders and the adapted encoders are frozen.
-Despite the cross-entropy $\mathcal{L}_{ce}$ adopted, we also adopt the proposed joint optimal transport loss $\mathcal{L}_{\mathrm{jot}}$ between joint zero-short representation and adapted representation to constrain the model.
 3.1 Preliminaries
@@
 An image $I$ is first divided into $M$ patches and then projected to latent space, which is denoted as $\bm{V}^{0}=\{\bm{v}_{1}^{0},\cdots,\bm{v}_{M}^{0}\}$.
-Afterward, a learnable class token $\bm{v}_{cls}^{0}$ is concatenated as the global representation, denoted as $[\bm{V}^{0},\bm{v}_{cls}^{0}]$.
-These tokens subsequently are fed into a sequence of transformer blocks:
-$[\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (1)
-where $l=1,\cdots,L_{I}$ and $L_{I}$ denote the number of vision transformer blocks in the image encoder of CLIP.
-Afterward, the final image representation in V-L space is obtained through an additional linear projection:
-$\bm{h}=f_{\texttt{VisionProj}}(\bm{v}_{cls}^{L_{I}}).$ (2)
-Text Encoding. Each text prompt is tokenized to $N$ tokens, denoted as $\bm{T}^{0}=\{\bm{t}_{1}^{0},\cdots,\bm{t}_{N}^{0}\}$. Likewise, these tokens then feed to text transformer blocks,
-$\bm{T}^{l}$ $=\texttt{Trans}^{l}_{\texttt{Text}}(\bm{T}^{l-1}),$ (3)
-where $l=1,\cdots,L_{T}$ and $L_{T}$ denote the number of transformer blocks in the text encoder of CLIP. Similarly, the final representation is obtained by the last token through a linear projection:
-$\bm{g}=f_{\texttt{TextProj}}(\bm{t}_{N}^{L_{T}}).$ (4)
+Afterward, a learnable class token $\bm{v}_{cls}^{0}$ is concatenated as the global representation, denoted as $\bm{V}^{0},\bm{v}_{cls}^{0}]$.
+These tokens are subsequently fed into a sequence of transformer blocks: $\bm{V}^{l},\bm{v}_{cls}^{l}]=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ Where $l=1,\cdots,L_{I}$ and $L_{I}$ denote the number of vision transformer blocks in the image encoder of CLIP.
+Afterward, the final image representation in V-L space is obtained through an additional linear projection: $\bm{h}=f_{\texttt{VisionProj}}(\bm{v}_{cls}^{L_{I}}).$
+Text Encoding.
+Each text prompt is tokenized to $N$ tokens, denoted as $\bm{T}^{0}=\{\bm{t}_{1}^{0},\cdots,\bm{t}_{N}^{0}\}$.
+Likewise, these tokens are then fed to text transformer blocks as $\bm{T}^{l}=\texttt{Trans}^{l}_{\texttt{Text}}(\bm{T}^{l-1}),$ Where $l=1,\cdots,L_{T}$ and $L_{T}$ denote the number of transformer blocks in the text encoder of CLIP.
+Similarly, the final representation is obtained by the last token through a linear projection as $\bm{g}=f_{\texttt{TextProj}}(\bm{t}_{N}^{L_{T}}).$
 Prediction.
@@
 The prediction can be performed as:
-$p(\bar{y}|I)=\frac{\text{exp}(sim(\bm{h},\bm{g}_{y})/\tau)}{\sum_{k=1}^{C}\text{exp}(sim(\bm{h},\bm{g}_{i}/\tau)},$ (5)
+$p(\bar{y}|I)=\frac{\text{exp}(sim(\bm{h},\bm{g}_{y})/\tau)}{\sum_{k=1}^{C}\text{exp}(sim(\bm{h},\bm{g}_{k}/\tau))},$ (1)
 where $sim(\cdot,\cdot),\tau$ denote the cosine similarity and a pre-defined temperature, respectively.
 Prompt Learning.
-Now, we will delve into how to perform prompt learning, which often appends some learnable tokens to text or vision branches as some transformer layers.
-Formally, we denote $\mathcal{S}_{\texttt{vision}}$ and $\mathcal{S}_{\texttt{text}}$ as the layers to integrate learnable prompt tokens for vision encoder and text encoder, respectively.
+Now, we will delve into how to perform prompt learning, which often appends some learnable tokens to text or vision branches as part of transformer layers.
+Formally, we denote $\mathcal{S}_{\texttt{vision}}$ and $\mathcal{S}_{\texttt{text}}$ as the layers to integrate learnable prompt tokens for the vision encoder and text encoder, respectively.
 We can denote the $K_{I}$ vision prompt tokens and $K_{T}$ text prompt tokens at $(l+1)$-th block as $\bm{P}^{l}=\{\bm{p}^{l}_{1},\cdots,\bm{p}^{l}_{K_{I}}\}$ and $\bm{Q}^{l}=\{\bm{q}^{l}_{1},\cdots,\bm{q}^{l}_{K_{T}}\}$, respectively.
 The prompting for vision encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{vision}}$) can be then presented as:
-[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{P}^{l-1},\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (6)
+$[~\underline{\hskip 8.5359pt}~,\bm{V}^{l},\bm{v}_{cls}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Vision}}([\bm{P}^{l-1},\bm{V}^{l-1},\bm{v}_{cls}^{l-1}]),$ (2)
 and for text encoder in $l$-th layer ($l\in\mathcal{S}_{\texttt{text}}$) can be then presented as:
-[¯$[~{}\underline{\hskip 8.5359pt}~{},\bm{T}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Text}}([\bm{Q}^{l-1},\bm{T}^{l-1}]).$ (7)
+$[~\underline{\hskip 8.5359pt}~,\bm{T}^{l}]$ $=\texttt{Trans}^{l}_{\texttt{Text}}([\bm{Q}^{l-1},\bm{T}^{l-1}]).$ (3)
 It is worth mentioning that this is known as Independent Vision-Language Prompting (IVLP), if $\bm{P}^{l}$ and $\bm{Q}^{l}$ are defined independently.
@@
 In order to adapt classification on a downstream dataset, we can simply update the prompt tokens via cross-entropy loss $\mathcal{L}_{CE}$ between predictions (Eq.
-5) and ground truth labels.
+1) and ground truth labels.
 3.2 Vision-Text Joint Optimal Transport
@@
 During training, this framework involves a CLIP with learnable prompts, i.e., the adapted model and a frozen pre-trained CLIP, while during inference, we only need the adapted CLIP.
-The overview of this framework is shown in Fig.
-2.
-A key principle in CLIP is that an instance has representations from both the text domain and the image domain.
-Therefore, we should constrain them simultaneously for each instance.
-Suppose we have a dataset $\mathcal{D}=\{(I_{i},y_{i})\}_{i=1}^{n}$, where $I_{i},y_{i}$, and $n$ denote the $i$th input image, its corresponding label, and number of samples, respectively.
+As discussed in Sec.
+2, constraining each modality separately can result in uneven constraint strength due to inherent imbalances.
+To address this, we emphasize a key principle of CLIP: each instance is jointly represented in both the text and image domains.
+Consequently, the two modalities should be constrained simultaneously at the instance level.
+Suppose we have a dataset $\mathcal{D}=\{(I_{i},y_{i})\}_{i=1}^{n}$, where $I_{i},y_{i}$, and $n$ denote the $i$th input image, its corresponding label, and the number of samples, respectively.
 For each instance $i\in[n]$, the vision representation is obtained by the vision encoder as $\bm{h}^{i}$.
@@
 We denote it as $\bm{X}_{\texttt{zs}}=\{\bm{x}_{\texttt{zs}}^{i}\}_{i=1}^{n}$.
-It is worth mentioning that $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{i}$ are obtained from the same input instance $(I_{i},y_{i})$, while with different CLIP models (i.e.
-one keep updating and one frozen).
-Our proposed regularization term is presented as: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the proposed constraint lies in Optimal Transport [23] is defined as:
-$\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})=\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),$ (8)
+It is worth mentioning that $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{i}$ are obtained from the same input instance $(I_{i},y_{i})$, while with different CLIP models (i.e., one keeps updating and one is frozen).
+Our proposed regularization term is presented as follows: Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the proposed constraint lies in Optimal Transport [36] is defined as:
+$\vskip-7.22743pt\mathcal{L}_{\mathrm{jot}}(\bm{X},\bm{X}_{\texttt{zs}})=\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),\vskip-7.22743pt$ (4)
 where $c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})$ denotes a cost function (e.g.
-linear and quadratic cost) between $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{j}$ $\bm{\gamma}$ denotes a transport map, $\mathcal{U}(\bm{a},\bm{b})$ denotes set of all valid transport maps,
-$\mathcal{U}(\bm{a},\bm{b})$ =$=\left\{\bm{\gamma}\in\mathbb{R}^{n\times n}_{+}\ \middle|\ \bm{\gamma}\mathbf{1}_{n}=\bm{a},\ \bm{\gamma}^{\top}\mathbf{1}_{n}=\bm{b}\right\},$ (9)
+linear and quadratic cost) between $\bm{x}^{i}$ and $\bm{x}_{\texttt{zs}}^{j}$ .
+$\bm{\gamma}$ denotes a transport map, $\mathcal{U}(\bm{a},\bm{b})$ denotes set of all valid transport maps,
+$\mathcal{U}(\bm{a},\bm{b})=\left\{\bm{\gamma}\in\mathbb{R}^{n\times n}_{+}\ \middle|\ \bm{\gamma}\mathbf{1}_{n}=\bm{a},\ \bm{\gamma}^{\top}\mathbf{1}_{n}=\bm{b}\right\},$ (5)
 $\bm{a}=\frac{1}{n}\mathbf{1}_{n},\quad\bm{b}=\frac{1}{n}\mathbf{1}_{n},$
@@
 Here, we focus more on how using OT-loss can potentially enlarge the feasible parameter space, which potentially results in easier optimization and better convergence.
-We first give a general definition of the loss of point-wise constraint: {definition} Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the point-wise constraint is defined as:
+We first give a general definition of the loss of point-wise constraint.
+Given $\bm{X}$ and $\bm{X}_{\texttt{zs}}$, the point-wise constraint is defined as:
 $\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})=\frac{1}{n}\sum_{i=1}^{n}c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{i}).$ (6)
-We observe that it can be presented in a similar form of optimal transport: {lemma} An equivalent form of Definition 3.3 is presented as:
+We observe that it can be presented in a similar form of optimal transport:
 $\mathcal{L}_{\mathrm{pw}}(\bm{X},\bm{X}_{\texttt{zs}})=\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j}),$ (7)
@@
 Proof.
-We first note that, clearly, according to Definition 3.2, $\hat{\gamma}\in\mathcal{U}(\bm{a},\bm{b})$ in Lemma 3.3 is a valid transport map. Accordingly,
+We first note that, clearly, according to Eq. 5, $\hat{\gamma}\in\mathcal{U}(\bm{a},\bm{b})$ in Eq. 7 is a valid transport map. Accordingly,
 $\underbrace{\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})}_{\mathcal{L}_{\mathrm{pw}}}\geq\underbrace{\min_{\bm{\gamma}\in\mathcal{U}(\bm{a},\bm{b})}\sum_{i=1}^{n}\sum_{j=1}^{n}\bm{\gamma}_{ij}\,c(\bm{x}^{i},\bm{x}_{\texttt{zs}}^{j})}_{\mathcal{L}_{\mathrm{jot}}}.$ (8)
@@
 $\mathcal{P}=\{\bm{P}^{l}\}_{l\in\mathcal{S}_{\texttt{vision}}},\mathcal{Q}=\{\bm{Q}^{l}\}_{l\in\mathcal{S}_{\texttt{text}}}$ denote all learnable tokens from different layers.
-$\lambda\geq 0$ is a tuning hyper-parameter.
+$\lambda\geq 0$ is a tuning hyperparameter.
 This objective can be considered as the generalized Lagrange function of the following optimization problem:
-$\min_{\mathcal{P},\mathcal{Q}}\quad\mathcal{L}_{ce}$ (14)
-$s.t.~{}R(\bm{X},\bm{X}_{\texttt{zs}})-\epsilon^{2}\leq 0,$
+$\min_{\mathcal{P},\mathcal{Q}}\quad\mathcal{L}_{ce}\quad s.t.~R(\bm{X},\bm{X}_{\texttt{zs}})-\epsilon^{2}\leq 0,$ (10)
 where $\epsilon^{2}$ is a certain constant.
@@
 Theorem 3.3 immediately suggests that using optimal transport constraints allows a broader feasible parameter domain of learnable tokens $(\mathcal{P},\mathcal{Q})$.
-A larger feasible domain often indicates there are potential more local minimal existing and provides more pathways for the optimizer to explore, which increases the chances of finding a better solution and alleviates a heavy hyperparameter search, i.e., a uniform hyperparameter can achieve reasonable performance across different datasets.
+A larger feasible domain often indicates there are potentially more local minima existing and provides more pathways for the optimizer to explore, which increases the chances of finding a better solution and alleviates a heavy hyperparameter search, i.e., a uniform hyperparameter can achieve reasonable performance across different datasets.
 3.4 Training and Inference
@@
 It is not intended for direct comparison but is included here for reference.
+Wilcoxon signed-rank test is shown in Appendix B.
 (a) Average
@@
 For Domain Generalization, we again train on ImageNet [5] and evaluate on four out-of-distribution datasets.
+Please refer to Appendix A for more details.
 Table 2: Performance of PromptOT on cross-dataset evaluation and its comparison to existing methods.
@@
 It is worth mentioning that a portion of the methods are only designed for partial tasks, while our method does not have any obstacles to implementation for all tasks.
-For the sake of a fair comparison and proof of concept, we will skip comparing methods that introduce substantial additional knowledge (e.g., LLMs or larger CLIP models for knowledge distillation) in our main experiments, such as those proposed in [22, 13, 32, 24].
+For the sake of a fair comparison and proof of concept, we will skip comparing methods that introduce substantial additional knowledge (e.g., LLMs or extra larger CLIP models for knowledge distillation) in our main experiments, such as those proposed in [34, 21, 46, 38].
 Implementation details.
@@
 Baseline results are taken from their respective papers.
-For more detail, refer to Appendix A.
+For more details, refer to Appendix A.
 4.2 Main Results
@@
 Cross-Dataset Evaluation. Table 2 shows that our method attains an average accuracy of 66.52% across 11 datasets, outperforming all baselines and exceeding PromptSRC by 0.62%.
-Domain Generalization. As shown in Table 3, we observe that our method achieves the highest average accuracy on four out-of-distribution datasets.
 Table 3: Performance on domain generalization. These baseline models are trained on ImageNet and evaluated on four out-of-distribution ImageNet datasets.
@@
 Prompt-OT 64.35 49.40 51.63 77.40 60.70
-Our method demonstrates robust performance across three key evaluations.
+Domain Generalization. As shown in Table 3, we observe that our method achieves the highest average accuracy on four out-of-distribution datasets.
+Overall, our method demonstrates robust performance across three key evaluations.
 In the base-to-novel generalization setup, it substantially outperforms zero-shot CLIP and surpasses the previous state-of-the-art in base accuracy, novel accuracy, and harmonic mean.
-In cross-dataset evaluation, it achieves the highest average accuracy over 11 diverse datasets, again surpassing all baselines.
-Finally, in domain generalization, our method obtains the best overall accuracy without relying on prompt ensembles or text augmentation, underscoring both its effectiveness and simplicity.
+In cross-dataset evaluation and in domain generalization, our method obtains the best overall accuracy without relying on prompt ensembles or text augmentation, underscoring both its effectiveness and simplicity.
 5 Analysis
@@
 All experiments will be conducted consistently with the setups of base-to-novel generalization.
-For fairness, we exclude all other components, and only apply regularization on the IVLP.
+For fairness, we exclude all other components and only apply regularization on the IVLP.
 Ours VS Point-wise Constraints.
 To illustrate the superiority of the proposed method over point-wise constraints, we evaluate it against several baselines: (i) naive $L_{2}$ constraints between $\bm{X}$ and $\bm{X}_{\texttt{zs}}$ with the same regularization strength $\lambda$, (ii) Adaptor-Cos [34], which introduces adapters at the end of the adapted encoders and employs cosine similarity as the point-wise constraint between representations, (iii) SRC [17], which applies $L_{1}$ losses for vision and text representations separately and a KL-divergence between prediction probabilities for each instance.
-(iv) Separate OT (see Appendix A), which applies OT-loss for vision and text representations separately.
-For a fair comparison, we exclude all additional augmentations, such as ensemble techniques and LLM-based augmentation.
-Our primary observation is that all these methods achieve a better balance between novel and base tasks than the model without any constraints, with a minimum improvement of 1.83% in harmonic mean accuracy.
-More importantly, the OT-based methods outperform all recent approaches based on point-wise constraints.
-Even without joint representation, applying OT for two domains separately can outperform all other point-wise methods.
-Additionally, compared to SRC, the previous state-of-the-art method, we achieved gains of 0.4%, 1.35%, and 0.93% in base task accuracy, novel task accuracy, and harmonic mean accuracy, respectively.
-Compared to $L_{2}$ constraints, we observe that with the same regularization strength $\lambda$, $L_{2}$ constraints, while preserving good generalizability to novel tasks, fail to yield any improvements on the base task.
-This suggests that $L_{2}$ constraints may be overly restrictive for model adaptation.
-In contrast, our proposed approach effectively balances performance across both the base and novel tasks, demonstrating that modeling cross-instance correlation while allowing the model greater flexibility in updating leads to an optimal trade-off, as expected in Fig.
-1.
-Table 4: Comparison of performance with point-wise constraints.
-Reg. Method Base Novel HM
+Table 4: Comparison of performance with point-wise constraints. Reported $p$-values are obtained from Wilcoxon signed-rank tests comparing the baseline and our method.
+Reg. Method Base Novel HM $p<0.05$
 w/o Constraints 84.20 71.21 77.16
-Point-wise $L_{2}$ between $\bm{X},\bm{X}_{\texttt{zs}}$ 84.20 74.68 79.16
-Adpator-Cos
-From CoPormpt [22]
+Point-wise
+$L_{2}$ between $\bm{X},\bm{X}_{\texttt{zs}}$ 84.20 74.68 79.16
+Adaptor-Cos
+From CoPrompt [34]
 83.12 75.25 78.99
@@
 Separate OT 84.50 75.48 79.74
-Our 84.81 76.25 80.30
+Ours 84.81 76.25 80.30
+(iv) Separate OT (see Appendix A), which applies OT-loss for vision and text representations separately.
+For a fair comparison, we exclude all additional augmentations, such as ensemble techniques and LLM-based augmentation.
+Table 4 reports both the performance metrics and the corresponding statistical test results.
+Our primary observation is that all these methods achieve a better balance between novel and base tasks than the model without any constraints, with a minimum improvement of 1.83% in harmonic mean accuracy.
+More importantly, the OT-based methods outperform all recent approaches based on point-wise constraints.
+Even without joint representation, applying OT for two domains separately can outperform all other point-wise methods.
+Additionally, compared to SRC, the previous state-of-the-art method, our proposed method achieved statistically significant gains of 0.4%, 1.35%, and 0.93% in base task accuracy, novel task accuracy, and harmonic mean accuracy, respectively.
+Compared to $L_{2}$ constraints, we observe that with the same regularization strength $\lambda$, $L_{2}$ constraints, while preserving good generalizability to novel tasks, fail to yield any improvements on the base task.
+This suggests that $L_{2}$ constraints may be overly restrictive for model adaptation.
+In contrast, our proposed approach effectively balances performance across both the base and novel tasks, demonstrating that modeling cross-instance correlation while allowing the model greater flexibility in updating leads to an optimal trade-off, as expected in Fig.
+1.
 Effectiveness of Joint Representation.
@@
 Specifically, we examine four different strategies: (i) imposing the constraint solely on the text representation, (ii) imposing the constraint solely on the vision representation, (iii) applying constraints separately to text and vision representation distributions (i.e., Separate OT), and (iv) our proposed joint representation, which considers an instance includes both vision and text representations.
-As shown in Table 5, applying constraints to both modalities yields a significantly greater improvement (2.58% and 3.14% in HM) compared to imposing the constraint on a single modality (0.62% and 0.96% in HM).
-Moreover, our proposed joint representation obtain more gain than applying constraints on vision and text distributions separately.
-Our method leads to the highest increase in HM (3.14%), suggesting that it provides the most effective balance between minimizing errors in both novel and base classes.
 Table 5: Ablation study on different modalities for imposing constraints.
@@
 Joint 84.81 76.25 80.30
+As shown in Table 5, applying constraints to both modalities yields a significantly greater improvement (2.58% and 3.14% in HM) compared to imposing the constraint on a single modality (0.62% and 0.96% in HM).
+Moreover, our proposed joint representation yields greater improvements compared to applying constraints separately on vision and text distributions, which potentially results in unbalanced constraint strength.
+Our method leads to the highest increase in HM (3.14%), suggesting that it provides the most effective balance between minimizing errors in both novel and base classes.
+Figure 2: The effectiveness of $\lambda$ on base-to-novel generalization tasks averaged over 11 datasets.
 Sensitivity of $\lambda$.
@@
 As shown in Fig.
-3, as $\lambda$ increases, the constraints become stronger, leading to a decline in adaptation to the base class.
+2, as $\lambda$ increases, the constraints become stronger, leading to a decline in adaptation to the base class.
 Meanwhile, the accuracy on the novel class initially increases before decreasing, indicating that the zero-shot weights may not be optimal for both novel and base tasks.
@@
 Empirically, we find that $\lambda=10$ achieves a good balance between generalization and adaptation.
-Figure 3: The effectiveness of $\lambda$ on base-to-novel generalization tasks averaged over 11 datasets.
+Efficiency.
+Our method introduces only a minor computational overhead during training, as shown in Table 6, incurring approximately a 1.8% increase in runtime with negligible additional GPU usage.
+Table 6: Comparison of Computational Cost.
+Runtime GPU
+w/o Constraint 0.0385s/step 3948MB
+Ours 0.0392s/step 3950MB
 6 Conclusion
@@
 The respective epochs are fixed across all datasets.
+All experiments are run on a node of the cluster with one V100 16GB NVIDIA GPU.
 Following a deep prompting strategy, we apply prompts to all transformer blocks for the first benchmark and the first three transformer blocks for the remaining benchmarks, and we use four text prompts and four vision prompts.
@@
 The accuracy is reported based on the evaluation of base classes and novel classes, respectively.
-A.1 Separate OT
+The overview of this framework is shown in Fig.
+S3.
+Figure S3: The overview of our proposed framework.
+Only the prompt tokens are trainable, and the rest of the weights in both the zero-shot encoders and the adapted encoders are frozen.
+Despite the cross-entropy $\mathcal{L}_{ce}$ adopted, we also adopt the proposed joint optimal transport loss $\mathcal{L}_{\mathrm{jot}}$ between joint zero-short representation and adapted representation to constrain the model.
+A.1 Datasets
+We used the following datasets in our experiments:
+Exps. 1 and 2. ImageNet [5], Caltech101 [7], OxfordPets [29], StanfordCars [20], Flowers102 [28], Food101 [1], and FGVCAircraft [24], SUN397 [39], UCF101 [35], DTD [3], and EuroSAT [12].
+Exp. 3. ImageNet [5] as a source dataset and use ImageNet-A [14], ImageNet-R [13], ImageNet-Sketch [37] and ImageNetV2 [32].
+A.2 Separate OT
 The way of applying separate OT is similar to SRC loss.
@@
 $\mathcal{L}_{\texttt{SOT}}=\mathcal{L}_{\texttt{OT}}(\bm{H},\bm{H}_{zs})+\mathcal{L}_{\texttt{OT}}(\bm{G},\bm{G}_{zs}).$ (13)
-A.2 Reproducibility
+A.3 Reproducibility
 The code and model weights will be publicly released upon acceptance. In the supplementary zip file, we provide the key code for the configuration and implementation of our proposed method.
 Our code is developed based on https://github.com/muzairkhattak/PromptSRC.
+Appendix B Wilcoxon signed-rank test
+We perform a Wilcoxon signed-rank test [4] to compare our method with PromptSRC, the previous state-of-the-art, using harmonic mean accuracy as the evaluation metric.
+As shown in Table S7, our method achieves statistically significant improvements over PromptSRC (p-value ¡ 0.05).
+Table S7: Wilcoxon signed-rank test on Exp 1.
+PromptSRC VS Prompt-OT
+P-value 0.016
+Appendix C Limitation
+While our proposed PromptOT framework demonstrates strong performance across various benchmarks and introduces theoretically grounded improvements over prior prompt learning approaches, we acknowledge certain limitations that open avenues for future exploration:
+First, the OT-based regularization introduces additional computational overhead compared to simpler constraints (e.g., point-wise L2 losses).
+Although we adopt mini-batch OT solvers to keep training efficient, the method still requires access to additional computational resources.
+Second, while our approach avoids relying on external augmentations or auxiliary models (e.g., large LLMs), this also means it may not fully leverage certain recent advances in generative or data-enhancement techniques.
+Integrating such components in a resource-efficient manner could further boost performance.